OpenStudy (tylermckinney16):

Ill give metals plus fan you so please help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (tylermckinney16):

Please help!!!!!!!!! question is here :)

OpenStudy (mathstudent55):

metals: lead, iron, aluminum, gold medals: awards

OpenStudy (mathstudent55):

Are you learning trig functions or the ratios of the sides of a 30-60-90 triangle?

OpenStudy (tylermckinney16):

yeah

OpenStudy (tylermckinney16):

@mathstudent55

OpenStudy (mathstudent55):

If you use trig functions, then look at your figure. In your triangle, which side is the hypotenuse?

OpenStudy (tylermckinney16):

h?

OpenStudy (mathstudent55):

Correct. A right triangle has a hypotenuse, in our case h, and two legs. The legs are called adjacent leg and opposite leg. They are adjacent and opposite with respect to one of the acute angles of the triangle. We are interested in the 30-deg angle. With respect to the 30-deg angle, is side 10 adjacent or opposite?

OpenStudy (tylermckinney16):

opposite?

OpenStudy (mathstudent55):

Correct. Opposite means across from. Side 10 is opposite the 30-degree angle. Now we see what we have. We have a 30 degree angle. We know the opposite leg is 10. We want the hypotenuse. Which trig function relates the hypotenuse to the opposite leg? Do you know SOHCAHTOA?

OpenStudy (tylermckinney16):

No i dont lol.

OpenStudy (mathstudent55):

\(\sin A = \dfrac{opp}{hyp} \) \(\cos A = \dfrac{adj}{hyp} \) \(\tan A = \dfrac{opp}{adj} \)

OpenStudy (mathstudent55):

Look at the sine, cosine, and tangent ratios above. Which one relates the opposite leg to the hypotenuse?

OpenStudy (tylermckinney16):

sinA

OpenStudy (mathstudent55):

\(\color{red}{\sin A} = \dfrac{\color{red}{opp}}{\color{red}{hyp}} \) \(\cos A = \dfrac{adj}{hyp} \) \(\tan A = \dfrac{opp}{adj} \)

OpenStudy (mathstudent55):

Correct. Now we just write an equation using the sine ratio, and we solve for the unknown.

OpenStudy (mathstudent55):

\(\sin 30^\circ = \dfrac{10}{h} \)

OpenStudy (mathstudent55):

Multiply both sides by h. Then divide both sides by sin 30 to find h.

OpenStudy (mathstudent55):

\(h \sin 30 ^\circ = 10\) \(h = \dfrac{10}{\sin 30^\circ} \)

OpenStudy (tylermckinney16):

is it 30?

OpenStudy (mathstudent55):

Now use a calculator to do the operation on the right side. Make sure the calculator is in degree mode.

OpenStudy (mathstudent55):

h is not 30

OpenStudy (mathstudent55):

\(\sin 30^\circ = 0.5\)

OpenStudy (mathstudent55):

What is 10/0.5 ?

OpenStudy (tylermckinney16):

20

OpenStudy (mathstudent55):

Good. Now that you have h, you know two sides of a right triangle. You can use the Pythagorean equation to find the third side. You can also use trig.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

\(a^2 + b^2 = c^2\)

OpenStudy (mathstudent55):

\(10^2 + g^2 = 20^2\) Solve for g.

OpenStudy (mathstudent55):

Sorry, but gtg.

OpenStudy (tylermckinney16):

is it 10?

OpenStudy (tylermckinney16):

The answer?

OpenStudy (anonymous):

h=20 and g=\[10\sqrt{3}\]

OpenStudy (tylermckinney16):

Thans and whats ?

OpenStudy (tylermckinney16):

^^^^ thats another question